VIBRONIC SPECTROSCOPY OF MgCH_{3} Ã ^{2}E←X̃ ^{2}A_{1} TRANSITION

\underline{ALDO P. SALZBERG}, BRIAN E. APPLEGATE, TERRY A. MILLER, Laser Spectroscopy Facility, Department of Chemistry, The Ohio State University, Columbus, OH 43210.

We report the laser excitation and dispersed fluorescence spectra of the MgCH_{3} Ã ^{2}E← X̃ ^{2}A_{1} transition. Improvements to the experiment have allowed the observation of important vibronic features not previously observed. The radical was produced in a free jet expansion from the simultaneous photolysis of Hg(CH_{3})_{2} and the laser vaporization of solid Mg. The use of Hg(CH_{3})_{2} as a precursor substancially improves production of the radical. A 0.1 cm^{−1}-resolution laser was used to obtain the partially rotationally resolved LIF spectra of the 6_{0}^{1}, and 2_{0}^{1} transitions. Analysis of these spectra, along with the previously reported^{a} 0_{0}^{0} band yield structural parameters which provide information about the radical’s Ã ^{2}E state. Besides the usual rotational constants, we have obtained information about the Jahn-Teller effect, which is dynamic and relatively weak. We have also obtained the values of the Coriolis coupling and the spin-orbit coupling constants. The fluorescence lifetime was found to be 60(7) ns for the vibrationless level of the Ã state and less than 15 ns for transitions involving vibrational excitation. The dispersed fluorescence spectra of the jet-cooled MgCH_{3} was analyzed, and all observed ^{2}A_{1} state vibrational features were assigned.

aR. Rubino, J. Williamson, and T. A. Miller, J. Chem. Phys. 103(14 (1995).